Sheri drove southeast from Albuquerque to Roswell and then south to Carlsbad. She averaged 75 mph along the route from Albuquerque to Roswell and averaged 70 mph from Roswell to Carlsbad. She covered the entire 220 miles in 3 hours. How long did she take to dri
5.ve from Albuquerque to Roswell?
A) 2 hours
B) 1 hour
C) 1.5 hours
D) 2.5 hours

Answer by checkley71(8403)

75X+70(3-X)=220
75X+210-70X=220
5X=220-210
5X=10
X=10/5
X=2 HOURS TO DRIVE FROM ALBUQUERQUE TO ROSWELL. ANSWER A)
3-2=1 HOURS FOR THE DRIVE FROM FROM ROSWELL TO CARLSBAD.
PROOF
75*2+70*1=220
150+70=220
220=220